# A survey of the monotonicity and non-contradiction of consensus methods and supertree methods

### v2

@@ 3.1.2 Non-contradiction
We will now turn our attention to general consensus methods; i.e., we will no longer just consider the established ones.
3.2 General consensus methods
3.2.1 Monotonicity
Theorem 3.4.
Proof.
Theorem 3.5.
Proof.
Example 3.6.
For the unrooted case, consider again Figure 1. Let $\mathcal{P}=(T_{2},T_{2},T_{3})$, i.e., $\mathcal{P}$ employs two copies of $T_{2}$ and one copy of $T_{3}$. For the rooted case, we introduce a root on the inner edges of $T_{2}$ and $T_{3}$, respectively. It can easily be seen that in both cases, we have $\mathtt{MR}(\mathcal{P})=T_{2}$, whereas $\mathtt{loose}(\mathcal{P})$ is the star tree (i.e., $T_{1}$ in Figure 1; in the rooted case the only inner vertex of $T_{1}$ is then the root). So in particular, we have $\mathtt{MR}(\mathcal{P})\not\preceq\mathtt{loose}(\mathcal{P})$. Since $\mathtt{MR}$ is unanimous (it is even regular, cf. [6]) and monotonic by Theorem 3.1, this shows that the conditions of Theorem 3.5 cannot be relaxed.
Proposition 3.7.
Suppose $\phi$ is a regular, monotonic consensus method. Let $\mathcal{P}=(T_{1},\ldots,T_{k})$ be a profile of $k$ rooted (or unrooted) phylogenetic $X$-trees, and let $c_{1}$ and $c_{2}$ be two clusters (or $\sigma_{1}$ and $\sigma_{2}$ be two splits) such that $c_{1}$ (or $\sigma_{1}$) is compatible with all trees in $\mathcal{P}$, while $c_{2}$ (or $\sigma_{2}$) is not compatible with $c_{1}$ (or $\sigma_{1}$, respectively). Then, $\phi(P)$ does not contain $c_{2}$ (or $\sigma_{2}$).
Proof.
As we will now show, the previous proposition implies that if there are two clusters (or splits) that are incompatible with one another, neither one of them can be contained in the tree generated by a regular and monotonic consensus method, even if both of them are compatible with all input trees.
Corollary 3.8.
Proof.
So just because two clusters (or splits) are incompatible with one another, they cannot be contained in the output of any regular and monotonic consensus method – even if both are compatible with all input trees. Again, note that by Theorem 3.1, this for instance applies to majority rule and strict consensus, i.e., to two of most frequently used consensus methods. Another direct consequence of Proposition 3.7 is the following corollary.
Corollary 3.9.
Proof.
A final consequence of Proposition 3.7 is the following corollary.
Corollary 3.10.
Proof.
3.2.2 Non-contradiction
Theorem 3.11.
Before we can prove Theorem 3.11, we need one more lemma.
Lemma 3.12.
Let $T$ be a phylogenetic $X$-tree with split set $\Sigma(T)$ (or cluster set $\mathcal{C}(T)$ in the rooted case) and let $\sigma\not\in\Sigma(T)$ (or $c\not\in\mathcal{C}(T)$) be compatible with all splits in $\Sigma(T)$ (or all clusters in $\mathcal{C}(T)$, respectively). Then, there is a refinement $T^{\prime}$ of $T$ which is incompatible with $c$.
Proof.
We start with the rooted case. Let $T$, $\sigma$ and $c$ be as described in the theorem. We first consider the rooted case. Let $\widetilde{T}$ be the unique tree induced by $\mathcal{C}(T)\cup\{c\}$. Then $\widetilde{T}$ has an edge $e$ that $T$ does not have, namely the one inducing $c$ (note that $\widetilde{T}$ must exist due to compatibility). Contracting this edge turns $\widetilde{T}$ into $T$ and also leads to a vertex $v$, which is ancestral to all leaves of $c$. However, as $c$ is not a cluster of $T$, at least two subtrees $T_{1}$, $T_{2}$ descending from $v$ in $T$ contain only leaves of $c$. But as $v$ does not induce $c$, there must be another subtree $T_{3}$ of $v$ which does not contain any leaf from $c$. Let $X_{i}$ denote the leaf set of $T_{i}$ for $i=1,2,3$, respectively. Then, let $c^{\prime}=X_{1}\cup X_{3}$ and let $T^{\prime}$ be the unique tree induced by $\mathcal{C}(T)\cup\{c^{\prime}\}$. Then, $c$ and $c^{\prime}$ are not compatible by construction and $T^{\prime}$ is a refinement of $T$. This completes the proof for the rooted case.
We are now finally in a position to prove Theorem 3.11.
Proof of Theorem 3.11.
Assume the theorem is false, i.e., assume that there are a monotonic and non-contradictory consensus method and a profile $\mathcal{P}=(T_{1},\ldots,T_{k})$ such that there is a cluster $c$ (or a split $\sigma$, respectively) in $\phi(\mathcal{P})$ not displayed by any $T_{i}$ (for $i=1,\ldots,k$). Then, as $\phi$ is non-contradictory, $c$ (or $\sigma$) is compatible with at least one tree $T_{i}$ (for $i\in\{1,\ldots,k$}). By Lemma 3.12 there is a refinement $T_{i}^{\prime}$ of $T_{i}$ which is incompatible with $c$ (or $\sigma$). We replace $T_{i}$ by $T_{i}^{\prime}$ in $\mathcal{P}$, and we repeat this procedure as long as there are trees compatbile with $c$. We call the resulting profile $\mathcal{P}^{\prime}$. As $\phi$ is monotonic, we know that $\phi(\mathcal{P})\preceq\phi(\mathcal{P}^{\prime})$. Thus, as $c$ (or $\sigma$) is contained in $\phi(\mathcal{P})$, it is also contained in $\phi(\mathcal{P}^{\prime})$. However, as $c$ is by construction incompatible with all trees in $\mathcal{P}^{\prime}$, this contradicts the fact that $\phi$ is non-contradictory. This shows that the assumption was wrong and thus completes the proof. ∎
While we have already seen that the $\mathtt{MR}$ and $\mathtt{strict}$ consensus methods are regular, monotonic and non-contradictory (and all these are desirable properties of any consensus method, after all), whereas $\mathtt{loose}$ is not (as it lacks monotonicity), the question remains if this combination makes $\mathtt{strict}$ and $\mathtt{MR}$ unique or if there are other consensus methods – biologically justified or not – that also share all these properties. Indeed, Day and McMorris [8, Section 1.2] divide major axiomatic consensus theory results into several categories, and a uniqueness result would be an excellent example of a result of the form ”Consensus [method] $C$ is the unique rule $C$ if and only if $C$ has the desirable properties $X,Y,$ and $Z$. Unfortunately, this is not the case. The following theorem, which is the main result of this section, states that there are in fact infinitely many such consensus methods, so these properties are not unique to the mentioned ones.
Theorem 3.13.
Proof.
For the rooted case, consider the profile $\mathcal{P}=(T_{1}^{\prime},T_{2})$, where $T_{1}$ and $T_{2}$ are the two phylogenetic $X$-trees depicted in Figure 3 and $T_{1}^{\prime}$ is a refinement of $T_{1}$ (not necessarily a strict one, so we may have $T_{1}^{\prime}=T_{1}$), and where $X=\{1,\ldots,n\}$. For the unrooted case, consider the exact same profile, but suppress the root, i.e., replace the two edges incident to the degree-2 root vertex by a single edge.
Note that the exact structure of the trees depends on whether $n=2m+2$ or $n=2m+3$ for some $m\in\mathbb{N}$, i.e., whether $n$ is even or odd. Note that $T^{*}_{1\ldots m}$ and $T^{*}_{m+1\ldots 2m}$ are identical if leaf labels are disregarded, i.e., they correspond to the same treeshape, say $T^{*}$, but $T^{*}$ can be chosen to be any rooted binary tree (note that this implies that there are $WE(m)$ many choices for $T^{*}$, where $WE(m)$ denotes the $m^{th}$ Wedderburn Etherington number [24, Sequence A001190]).
We now define a consensus method $\phi$ as follows: If $\widetilde{\mathcal{P}}=\mathcal{P}$, i.e., if $\widetilde{\mathcal{P}}$ contains precisely $T_{1}^{\prime}$ and $T_{2}$, then we define $\phi(\widetilde{\mathcal{P}}):=T_{2}$. Let $\pi$ be a permutation on $X$. Let $\pi(T)$ denote a version of $T$ in which the leaves are permuted according to $\pi$, where $T$ is a phylogenetic $X$-tree. Then, we also define $\phi(\pi(T_{1}^{\prime}),\pi(T_{2})):=\pi(T_{2})$.
We now argue that $\phi$ is regular, monotonic and non-contradictory. Unanimity follows by the unanimity of $\mathtt{strict}$, and neutrality follows by $\mathtt{strict}$ for all profiles that are not $\widetilde{P}$, and for $\widetilde{P}$ it follows by our consideration of permutations $\pi$. To check anonymity, note that $T_{1}^{\prime}$ is either such that $T_{1}^{\prime}\preceq T_{2}$ or such that $T_{1}$ and $T_{2}$ have completely different tree shapes (in particular if $T_{1}$ gets refined in a way that does not form $T_{2}$, namely by combining leaf $12m+$ 1 with one of the copies of $T^{*}$ to form a new subtree). This shows that the only time when a permutation can modify $T_{1}^{\prime}$ to give $T_{2}$ and $T_{2}$ to give $T_{1}^{\prime}$ is when $T_{1}^{\prime}$ and $T_{2}$ are isomorphic. However, in this case we actually have by unanimity that $\phi$ returns precisely this tree, so for anonymity, there remains nothing to show. Thus, $\phi$ is regular.
To see that $\phi$ is actually monotonic, recall that $\mathtt{strict}$ is monotonic by Theorem 3.1, and if a profile $\widetilde{\mathcal{P}}$ consists of $\pi(T_{1}^{\prime})$ and $\pi(T_{2})$ for some permutation $\pi$, then $\phi$ returns $\pi(T_{2})$, whether $T_{1}^{\prime}$ equals $T_{1}$ or is a strictly refined version of it. Thus, in summary, $\phi$ is monotonic.
Last, we need to show that $\phi$ is non-contradictory. To see this, recall that by Proposition 3.3, $\mathtt{strict}$ is non-contradictory, and if a profile $\widetilde{\mathcal{P}}$ consists of $\pi(T_{1}^{\prime})$ and $\pi(T_{2})$ for some permutation $\pi$, then $\phi$ returns $\pi(T_{2})$. So in particular, all splits of $\phi((\pi(T_{1}^{\prime}),\pi(T_{2})))=\pi(T_{2})$ are not only compatible with at least one of the input trees, but even contained in it, as $\pi(T_{2})$ is contained in the input profile, which implies that $\phi$ is co-Pareto.
Figure 3: In the proof of Theorem 3.13, we have $\mathcal{P}=(T_{1}^{\prime},T_{2})$ on $n=2m+2$ or $n=2m+3$ leaves, respectively, where $T_{1}^{\prime}$ is a refinement of $T_{1}$ as depicted here. $T^{*}_{1\ldots m}$ and $T^{*}_{m+1\ldots 2m}$ are both trees with the same binary treeshape $T^{*}$, e.g., a caterpillar (i.e., a rooted binary tree with precisely one cherry, i.e., one pair of leaves sharing the same parent), but the leaves of $T^{*}_{1\ldots m}$ are bijectively labelled by $\{1,\ldots,m\}$ and the leaves of $T^{*}_{m+1\ldots 2m}$ are bijectively labelled by $\{m+1,\ldots,2m\}$. Note that the dashed parts of the trees only exist if $n$ is odd.
3.3 Results on MRP and MRC
@@
Consider profile $\mathcal{P}=(T_{1},T_{2})$ consisting of the two trees $T_{1}$ and $T_{2}$ from Figure 4. Moreover, consider the corresponding alignment of these two trees as given in Table 1. It can be easily seen that neither $T_{1}$ nor $T_{2}$ are MP trees for this alignment; in fact, the alignment has parsimony score 13 on tree $T_{3}$ from Figure 4, but 14 on both $T_{1}$ and $T_{2}$, respectively (cf. Table 1). Moreover, we performed an exhaustive search over all phylogenetic trees with seven leaves using the computer algebra system Mathematica [26] in order to verify that $T_{3}$ is even the unique MP tree for this alignment. However, $T_{3}$ contains the split $\sigma=1,2,3,4|5,6,7$, which is not contained in any of the input trees $T_{1}$ and $T_{2}$. As $T_{1}$ and $T_{2}$ are binary, this (by the Buneman theorem) necessarily means that $\sigma$ is incompatible with both trees, because no more split can be added to said trees. This completes the proof.
@@
In order to complement the results of [6], we briefly consider MRP and MRC in the light of adding more taxa (rather than more splits or clusters). This scenario is for instance relevant concerning the discovery of new species. The results stated in the following and the construction of the particular examples were strongly inspired by [11]. As stated above, both MRP and MRC are strictly speaking not consensus methods as their output may be a set of trees rather than a single tree. However, we will here focus on a profile $\mathcal{P}$ for which both MRP and MRC produce a unique tree, namely $\widetilde{\mathcal{P}}=(\widetilde{T_{1}},\widetilde{T_{1}},\widetilde{T_{2}},\widetilde{T_{2}},\widetilde{T_{2}},\widetilde{T_{3}})$, where $\widetilde{T_{1}},\widetilde{T_{2}},\widetilde{T_{3}}$ are as depicted in Figure 5. Now, we add taxon 5 to all trees in $\widetilde{\mathcal{P}}$ such that we get profile $\mathcal{P}=(T_{1},T_{1},T_{2},T_{3},T_{4},T_{5})$, where again $T_{1},\ldots,T_{5}$ are as depicted in Figure 5. Note that $T_{1}\setminus\{5\}=\widetilde{T_{1}}$, $T_{2}\setminus\{5\}=T_{3}\setminus\{5\}=T_{4}\setminus\{5\}=\widetilde{T_{2}}$, and $T_{5}\setminus\{5\}=\widetilde{T_{3}}$. Translating all non-trivial splits of $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ to characters leads to alignments $\mathcal{A}_{\mathcal{P}}$ and $\mathcal{A}_{\widetilde{\mathcal{P}}}$ as depicted in Figures 6 and 7.
@@
Figure 5: Profile $\widetilde{\mathcal{P}}$, which can be turned into profile $\mathcal{P}$ by adding information on taxon 5, leading to alignments $\mathcal{A}_{\widetilde{\mathcal{P}}}$ and $\mathcal{A}_{\mathcal{P}}$ as depicted in Figures 6 and 7, respectively. $T_{1}$ is the unique MRP and MRC tree for $\mathcal{P}$, while $\widetilde{T_{2}}$ is the unique MRP and MRC tree for $\widetilde{T}$.
@@
We used the computer algebra system Mathematica [26] to prove by exhaustive search that $T_{1}$ is both the unique MRP and MRC tree for $\mathcal{A}_{\mathcal{P}}$, and that $\widetilde{T_{2}}$ is both the unique MRP and MRC tree for $\mathcal{A}_{\widetilde{\mathcal{P}}}$. However, as $\widetilde{T_{2}}$ is not a subtree of $T_{1}$, this shows that even in the case where MRP and MRC agree and additionally lead to a unique tree, additional taxa can change the predetermined relationships of the species in question. Thus, MRP and MRC are not ‘future-proof’ in the sense of [6].

### v4

@@ 3.1.2 Non-contradiction
We will now turn our attention to general the supertree methods MRP and MRC for the cases when they can easily be used as consensus methods, namely when their output is a unique tree.
3.2 Results on the established supertree methods MRP and MRC used as consensus methods
@@
Similarly to the previous section, where we showed that while monotonicity cannot be generally guaranteed for $\mathtt{loose}$, $\varphi_{Adams}$ and $\varphi_{Aho}$, it does hold in the case where the finer input profile allows for a linear ordering (cf. Proposition 3.3, we want to consider this setting now for MRP and MRC.
Proposition 3.6.
Let $\mathcal{P}=\{T_{1},\ldots,T_{k}\}$ and let $\mathcal{P}^{\prime}=\{T_{1}^{\prime},\ldots,T_{k}^{\prime}\}$ be such that $T_{i}\preceq T_{i}^{\prime}$ for all $i\in\{1,\ldots,k\}$. Moreover, let $\mathcal{P}^{\prime}$ be such that it permits a linear ordering, i.e., for all $i,j\in\{1,\ldots,k\}$ we either have $T_{i}^{\prime}\preceq T_{j}^{\prime}$ or $T_{j}^{\prime}\preceq T_{i}^{\prime}$. Then we have: $MRC(\mathcal{P})$ and $MRC(\mathcal{P}^{\prime})$ are unique trees, and we have $MRC(\mathcal{P})\preceq MRC(\mathcal{P}^{\prime})$. Moreover, $MRC(\mathcal{P}^{\prime})$ is then also an MRP tree for $\mathcal{P}$ and $\mathcal{P}^{\prime}$, and $MRC(\mathcal{P})$ is also an MRP tree for $\mathcal{P}$. If additionally we have that $\mathcal{P}^{\prime}$ contains a binary tree $T$, then $MRP(\mathcal{P})=MRC(\mathcal{P}^{\prime})=T$.
Proof.
Let $\mathcal{P}$ and $\mathcal{P}^{\prime}$ be as described in the proposition. Then, as $\mathcal{P}^{\prime}$ allows for a linear ordering, we may assume without loss of generality that $T_{i}^{\prime}\preceq T_{i+1}^{\prime}$ for all $i=1,\ldots,k-1$. Note that this implies $T_{k}^{\prime}$ is the most refined tree in $\mathcal{P}^{\prime}$, and as the split set $\Sigma(\mathcal{P}^{\prime})$ of all splits in $\mathcal{P}^{\prime}$ contains only elements which are pairwise compatible (by $T_{i}^{\prime}\preceq T_{i+1}^{\prime}$), we know by the Buneman theorem that there is a unique tree containing all splits from $\Sigma(\mathcal{P}^{\prime})$. Necessarily, this tree must be $T_{k}^{\prime}$, which shows that $MRC(\mathcal{P}^{\prime})=T_{k}^{\prime}$, so $MRC(\mathcal{P}^{\prime})$ is indeed a unique tree. Moreover, as we have $T_{i}\preceq T_{i}^{\prime}\preceq T_{k}^{\prime}$ for all $i=1,\ldots,k$, we know that $\Sigma(\mathcal{P})\subseteq\Sigma(T_{k}^{\prime})$, which implies that all splits in $\Sigma(\mathcal{P})$ are compatible. Thus, again by the Buneman theorem, there is a unique tree $T$ with split set $\Sigma(\mathcal{P})$, which necessarily is the unique MRC tree of $\mathcal{P}$. Moreover, by $\Sigma(T)=\Sigma(\mathcal{P})\subseteq\Sigma(\mathcal{P}^{\prime})=\Sigma(MRC(\mathcal{P}^{\prime}))=\Sigma(T_{k}^{\prime})$ we conclude $MRC(\mathcal{P})\preceq MRC(\mathcal{P}^{\prime})$ as desired.
[…]
@@
[…]
@@
In order to complement the results of [9], we briefly consider MRP and MRC in the light of adding more taxa (rather than more splits or clusters). This scenario is for instance relevant concerning the discovery of new species. The results stated in the following and the construction of the particular examples were strongly inspired by [26]. As stated above, both MRP and MRC are strictly speaking not consensus methods as their output may be a set of trees rather than a single tree. However, we will here focus on a profile $\mathcal{P}$ for which both MRP and MRC produce a unique tree, namely $\widetilde{\mathcal{P}}=(\widetilde{T_{1}},\widetilde{T_{1}},\widetilde{T_{2}},\widetilde{T_{2}},\widetilde{T_{2}},\widetilde{T_{3}})$, where $\widetilde{T_{1}},\widetilde{T_{2}},\widetilde{T_{3}}$ are as depicted in Figure 4. Now, we add taxon 5 to all trees in $\widetilde{\mathcal{P}}$ such that we get profile $\mathcal{P}=(T_{1},T_{1},T_{2},T_{3},T_{4},T_{5})$, where again $T_{1},\ldots,T_{5}$ are as depicted in Figure 4. In the following, let the notation $T\setminus Y$ for a phylogenetic $X$-tree $T$ and a proper subset $Y$ of $X$ denote the phylogenetic $(X\setminus Y)$-tree that can be obtained from $T$ by deleting all leaves of $Y$ and suppressing all resulting degree-2 vertices as well as all newly occurring (i.e., unlabelled) leaves. Then, we have: $T_{1}\setminus\{5\}=\widetilde{T_{1}}$, $T_{2}\setminus\{5\}=T_{3}\setminus\{5\}=T_{4}\setminus\{5\}=\widetilde{T_{2}}$, and $T_{5}\setminus\{5\}=\widetilde{T_{3}}$. Translating all non-trivial splits of $\mathcal{P}$ and $\widetilde{\mathcal{P}}$ to characters leads to alignments $\mathcal{A}_{\mathcal{P}}$ and $\mathcal{A}_{\widetilde{\mathcal{P}}}$ as depicted in Figures 5 and 6.
@@
Figure 4: Profile $\widetilde{\mathcal{P}}$, which can be turned into profile $\mathcal{P}$ by adding information on taxon 5, leading to alignments $\mathcal{A}_{\widetilde{\mathcal{P}}}$ and $\mathcal{A}_{\mathcal{P}}$ as depicted in Figures 5 and 6, respectively. $T_{1}$ is the unique MRP and MRC tree for $\mathcal{P}$, while $\widetilde{T_{2}}$ is the unique MRP and MRC tree for $\widetilde{T}$.
We now argue that $T_{1}$ is both the unique MRP and MRC tree for $\mathcal{A}_{\mathcal{P}}$, and that $\widetilde{T_{2}}$ is both the unique MRP and MRC tree for $\mathcal{A}_{\widetilde{\mathcal{P}}}$. First note that resolving a tree more cannot increase the parsimony score, but it can potentially decrease it, i.e., for $T\preceq T^{\prime}$ and a binary character $f$, we have $l(T^{\prime},f)\leq l(T,f)$. This holds because if multiple change edges are incident to the same vertex, replacing this vertex by a new edge might make it possible to have a change on the new edge rather than on multiple incident edges. Thus, when hunting for a tree minimizing the parsimony score, it is sufficient to consider binary trees. Now, as every binary character on four or five taxa requires at most 2 changes on any tree (it can always be realized with changes on the two edges incident to leaves that are not in the majority state), for a given character and a given tree with $n\leq 5$, there are only two possibilities: in case of compatibility, the parsimony score is 1, otherwise it is 2. So there are no different degrees of incompatibility (higher parsimony scores) for certain characters than for others, which is why it is clear that MRP and MRC necessarily coincide in this example: MRP will simply return the binary tree whose splits correspond to the characters occurring in total most frequently in $\mathcal{A}_{\mathcal{P}}$ and $\mathcal{A}_{\widetilde{\mathcal{P}}}$, which is precisely the MRC tree. While finding the MRC set in $\mathcal{A}_{\mathcal{P}}$ is a bit more involved, finding it in $\mathcal{A}_{\widetilde{\mathcal{P}}}$ is simple, as for four taxa, there are only three possible non-trivial splits, all of which are pairwise incompatible. Two of these splits’ corresponding characters are present in $\mathcal{A}_{\widetilde{\mathcal{P}}}$, and due to their incompatibility, MRC will simply pick the one that occurs most frequently, which is $ACCA$, leading to tree $\widetilde{T}_{2}$. Similarly, MRC will choose $T_{1}$ for $\mathcal{A}_{\mathcal{P}}$.
We now briefly describe why $T_{1}$ is the unique MRC tree for $A_{\mathcal{P}}$. MRC requires finding a maximum compatible subset of characters displayed by the input trees. Namely, this is $3$ copies of $AAACC$, $2$ copies each of $AACCC$ and $ACCAC$, and $1$ copy each of $AACAC$, $ACCCA$, $ACCAA$ and $ACAAC$. Tree $T_{1}$ displays a subset of $5$ of these ($3$ copies of $AAACC$ and $2$ copies each of $AACCC$), so it suffices to show that this is the unique tree that displays a compatible subset of size $5$ or more. This will follow from several observations. Firstly, $ACCAC$ is incompatible with both $AAACC$ and $AACCC$, and secondly the pair $AAACC$ and $AACCC$ are not both compatible with any of the remaining characters. It follows that either our maximum subset consists of 3 copies of $AAACC$ and 2 copies of $AACCC$, or our subset contains all copies of at most one of $AAACC,AACCC$ and $ACCAC$ together with some subset of the remaining characters (of which there are one of each). Of these, $AAACC$ is only compatible with $ACCAA$ (a maximum size $4$ subset), $AACCC$ is only compatible with $AACAC$ (a maximum size $3$ subset) and while $ACCAC$ is compatible with $AACAC$, $ACCAA$ and $ACAAC$ (potentially a size $5$ subset), $AACAC$ and $ACCAA$ are incompatible, leading again to a maximum size $4$ subset. It follows that the unique maximal subset of size $5$ is 3 copies of $AAACC$ and 2 copies of $AACCC$, which is uniquely displayed by $T_{1}$.
For the reasons explained above, these trees are then also the unique MRP trees in this case.
@@
However, as $\widetilde{T_{2}}$ is not a subtree of $T_{1}$, this shows that even in the case where MRP and MRC agree and additionally lead to a unique tree, additional taxa can change the predetermined relationships of the species in question. Thus, MRP and MRC are not ‘future-proof’ in the sense of [9].